ACETYLENE DISPERSED FLUORESCENCE SPECTRA ABOVE THE ISOMERIZATION BARRIER

MATTHEW P. JACOBSON, MICHELLE S. SILVA, ZICHENG DUAN, \underline{ROBERT W. FIELD}, Department of Chemistry and George R. Harrison Spectroscopy Laboratory, MIT, 77 Mass. Ave., Cambridge, MA 02139.

Dispersed fluorescence spectra of the acetylene Ã → X̃ system have been recorded, and observed transitions in these spectra have been assigned which terminate on eigenstates with greater than 15,000 cm^{−1} of internal energy. Results for both ^{13}C_{2}H_{2} and ^{12}C_{2}H_{2} will be discussed, as well as implications for acetylene-vinylidene isomerization, which is expected to become feasible at 15,200 cm^{−1}.^{a}

aJ. F. Stanton and J. Gauss, J. Chem. Phys. 110, 1831 (1999).